I am trying to make sure my answer is correct.
-(-2+4x)-(3-4x)+5=-(-3+6x)+x+1
My answer is 0=x

Answer by whisper(27)

Let's start by getting rid of the parenthesis:
-(-2+4x)-(3-4x)+5=-(-3+6x)+x+1
2-4x-3+4x+5=3-6x+x+1
combine terms:
2-3+5-4x+4x=3+1-6x+x
4-0=4-5x
4=4-5x
Subtract 4 from each side:
4-4=4-4-5x
0=0-5x
0=-5x
Divide both sides by -5
0/-5=-5x/-5
0=1x
0=x
Please don't hesitate to ask if you have further questions!